OpenStudy (anonymous):

Find the sum of this series: 3^0+3^1+3^2+3^3...+3^100

OpenStudy (anonymous):

This is a geometric progression.

OpenStudy (callisto):

sum = 3(1-3^100)/(1-3) +1 = 7.73x10^47

OpenStudy (anonymous):

Plz explain

OpenStudy (anonymous):

Hmmm.....that formula can't be true for "r>1".

OpenStudy (anonymous):

Technically, \( 1.(3^{101} -1)/2\)

OpenStudy (anonymous):

How?

OpenStudy (callisto):

i think the formula works..

OpenStudy (anonymous):

Yeah its one of my answer choices

OpenStudy (anonymous):

But explain how you got it plz

OpenStudy (anonymous):

@Callisto: It works here but not in general. Try taking limit where \( n\to \infty \)

OpenStudy (anonymous):

Isnt limit 100?

OpenStudy (slaaibak):

s = 1 + x + x^2 + ... + x^n xsn = x + x^2 + x^3 + ... + x^(n+1) xs - s = x^(n+1) - 1 s = (x^(n+1) -1)/(x-1)

OpenStudy (callisto):

then that would be GS sum to infinity that would be sum = a(1-r)/r

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@Callisto - you keep using the formula that works for r<1.

OpenStudy (anonymous):

slaaibak has it right.

OpenStudy (anonymous):

Thanks everybody medals for everyone :D

OpenStudy (callisto):

sorry that should be a/(1-r)

OpenStudy (callisto):

for me r<1 or r>1 is the same, just changing 1-r^n to r^n-1 and change the base also.. that's means taking out -1 as the common factor lol

OpenStudy (anonymous):

@Callisto: YOu will see the diff when you take the limit

OpenStudy (amistre64):

\[\lim_{n\to inf}\frac{1-r^n}{1-r}=\frac{-inf}{1-r}=-\frac{inf}{1-r}\] \[\lim_{n\to inf}\frac{r^n-1}{r-1}=\frac{inf}{r-1}=\frac{inf}{-(1-r)}=-\frac{inf}{1-r}\] so were am I missing this?

OpenStudy (callisto):

read the second post also

OpenStudy (amistre64):

sum = 3(1-3^100)/(1-3) +1 = 7.73x10^47 this one?

OpenStudy (amistre64):

A = 3^0+3^1+3^2+3^3+...+3^100 -3A = -3^1-3^2-3^3- ...-3^100-3^101 ---------------------------------------- (1-3)A = 3^0 -3^101 A = 3^0-3^101 --------- 1-3